# Multi-Agent Collaborative Data Selection for Efficient LLM Pretraining

### v1

@@ 3.2 Multi-agent Collaborative Data Selection Framework
Now the reward signal is actually came from multiple agents, the optimization goal in Equation 2 becomes maximizing the expectation of collaborative agents $\mathbb{E}_{\{\mathcal{A}_{1},...,\mathcal{A}_{n}\}}\left[R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{D}_{\text{ref}})\right]$. In our current implementation, we include three agents, which are topic agents, quality agents and domain agents. They are aiming to maximize the rewards from topic, quality and domain perspective respectively. In the following sections, we will detail how we initialize and update a single agent (§3.3), and how we update the agent console for multi-agent collaboration (§3.4).
@@
3.3 Single Agent Initialization and Update
Agent initialization.
As defined in Definition 1, for a particular agent, we have to maintain its memory $\mathcal{H}_{\mathcal{A}}$ throughout the training process. Before training process begin, we label the whole training dataset $\mathcal{D}$ offline and store the labled information to the memory of corresponding agents. Specifically, for each data point $x_{i}\in\mathcal{D}$, i.e., a single document in our settings, we first get the quality, topic and domain label using scorer and classifier.
For quality agent, we adopt the FineWeb-Edu quality scorer (Lozhkov et al., 2024), which is fine-tuned as a BERT-like regression model Merrick et al. (2024) using Llama3-70B-Instruct annotated 500k examples. This will give out a successive quality score $\text{Quality}(x_{i})\in\mathcal{R}^{[0,5]}$ with higher score represent higher quality. We then map this score into five quality intervals $\{I_{j}\}_{j=1}^{5}$, as
$\text{Quality}(x_{i})\mapsto I_{j}=\begin{cases}[j-1,j),&\text{if }\text{Quality}(x_{i})\in[j-1,j)\quad(j=1,2,3,4)\\ [4,5],&\text{if }\text{Quality}(x_{i})\in[4,5]\quad(j=5)\end{cases}$ (3)
We store the quality interval corresponding to each data point in the quality agent’s memory.
For the domain agent, we use the document’s meta-information, label the data with domain information and save this into the domain agent’s memory, where the domain $\text{Domain}(x_{i})$ belongs to one of ArXiv, Book, Wikipedia, CommonCrawl, GitHub, StackExchange, C4.
For the topic agent, due to the absence of a suitable pretrained model for topic classification and labeling, we designed a classification schema using 1.44 billion documents collected by the Common Crawl project (Project, 2007) and fine-tuned a BERT-like regression model on 500k GPT-4o annotated samples, the overall pipeline is depicted in Figure 3. Further details on the topic classification approach and BERT model training are provided in §A.1. Using this topic classifier, we categorize each document into one of 13 topics: Activity, Education, Entertainment, Finance, Health, Business and Industrial, Infrastructure, Literature and Art, Nature, Others, Law and Government, Networking, Technology, and store the topic information in the topic agent’s memory.
We initialize the weight of the topic agent and domain agent following the RegMix Liu et al. (2024) framework. Unlike the original RegMix, which only considers mixing data based on domain labels, we examine data mixing weights based on domain as well as the topic labels. We initialize our quality agent similar to the data selection decision of QuRating Wettig et al. (2024) and FineWeb-Edu Lozhkov et al. (2024). Further details can be found in §A.2.3. The initial weights for each agent are stored in their respective memory.
During the training phase, we leverage the current model to adjust the weight of each agent. As depicted in Figure 2, at the data selection stage, each agent performs several actions to update its memory and inform decision-making. Take domain agent as an example, it takes three-step action during data selection stage: (1) Sample pretraining data points from the pretraining data pool, distributing them uniformly across each domain; (2) Call the current model to assess the reward of each data point and gather feedback; (3) Update the memory of domain weights based on gathered feedback and adjusts the score for each data point by incorporating prior knowledge from the offline labeling process. This process is similarly followed by the quality agent and the topic agent.
Agent update.
After sampling data uniformly from agent search space, each agent updates its internal weights using local information based on the sampled data. For example, for domain agent, it calculates the average influence of each domain. For agent $\mathcal{A}\in\{\mathcal{A}_{\text{Quality}},\mathcal{A}_{\text{Domain}},\mathcal{A}_{\text{Topic}}\}$, it updates its internal weights by calculating the overall rewards sampled from each interval as:
$\overline{R}_{\mathcal{A}}^{j}=\frac{1}{|D_{\mathcal{A}}^{j}|}\sum_{x_{i}\in D_{\mathcal{A}}^{j}}\mathcal{R}(x_{i}\mid\mathcal{M},\mathcal{D}_{\text{ref}}),$ (4)
where $D_{\mathcal{A}}^{j}$ represents the sample set of the $j$-th subcategory under agent $\mathcal{A}$, e.g. Wikipedia for domain agent. And $x_{i}$ is a sample within this sample set. Then, the policy gradient method is used to update the weight for each subcategory $w_{\mathcal{A}}^{j}$:
$w_{\mathcal{A}}^{j}\leftarrow w_{\mathcal{A}}^{j}+\eta_{A}\cdot\overline{R}_{\mathcal{A}}^{j},$ (5)
where $\eta_{\mathcal{A}}$ is the learning rate for agent $\mathcal{A}$. The overall updated weight of agent $\mathcal{A}$ is an vector in $n_{\mathcal{A}}$ dimension, $\mathbf{w}_{\mathcal{A}}=[w_{\mathcal{A}}^{1},...,w_{\mathcal{A}}^{n_{\mathcal{A}}}]$, where $n_{\mathcal{A}}$ is the number of total subcategory within the space of agent $\mathcal{A}$. Utilizing the prior memory stored by each agent, it can give out a final score for each data point as $\mathcal{S}_{\mathcal{A}}(x_{i})=w_{\mathcal{A}}^{j}$, where $j$ is the subcategory that $x_{i}$ belongs to.
Calculating the rewards.
For each sampled data point, we approximate rewards using influence functions Engstrom et al. (2024); Yu et al. (2024). The influence function value is computed to measure the impact of each sample on the model’s performance. The formula for the influence function is:
$\mathcal{R}(x_{i}\mid\mathcal{M},\mathcal{D}_{\text{ref}})$ $=\hat{\mathbb{E}}_{x\sim\mathcal{D}_{\text{ref}}}\left[-\ell(x;\mathcal{O}(\mathcal{M},\mathcal{D}_{k}))\right],$ (6)
$=-\nabla_{\mathcal{M}}\mathcal{L}(\mathcal{D}_{\text{ref}}\mid\mathcal{M})^{\top}H^{-1}_{\mathcal{M}}\nabla_{\mathcal{M}}\mathcal{L}(x_{i}\mid{\mathcal{M}}),$
where $H_{\mathcal{M}}=\frac{1}{n}\sum_{j=1}^{n}\nabla^{2}_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}(x_{i}\mid{\mathcal{M}})$ is the Hessian and its positive definite. Details of calculating influence functions for pretraining data point can be found in §A.4.
3.4 Multi-agent Collaboration
Algorithm 1 Multi-agent collaborative data selection for LLM pretraining

### v3

@@ 3.2 Multi-Actor Data Selection Framework
Now the reward signal is actually came from multiple actors, the optimization goal in Equation 2 becomes maximizing the expectation of collaborative actors. In our current implementation, we include three actors, which are topic actors, quality actors and domain actors. They are aiming to maximize the rewards from topic, quality and domain perspective respectively. In the following sections, we will detail how we initialize and update a single actor (§3.3), and how we update the actor console for multi-actor collaboration (§3.4).
3.3 Single Actor Initialization and Update
Actor initialization.
As defined in Definition 1, for a particular actor, we have to maintain its memory $\mathcal{H}_{\mathcal{A}}$ throughout the training process. Before training process begin, we label the whole training dataset $\mathcal{D}$ offline and store the labeled information to the memory of corresponding actors. Specifically, for each data point $x_{i}\in\mathcal{D}$, i.e., a single document in our settings, we first get the quality, topic and domain label using scorer and classifier.
We initialize the weight of the topic actor and domain actor following the RegMix Liu et al. (2024) framework. Unlike the original RegMix, which only considers mixing data based on domain labels, we examine data mixing weights based on domain as well as the topic labels. We initialize our quality actor similar to the data selection decision of QuRating Wettig et al. (2024) and FineWeb-Edu Lozhkov et al. (2024). Further details can be found in §A.7.3. The initial weights for each actor are stored in their respective memory.
During the training phase, we leverage the current model to adjust the weight of each actor. As depicted in Figure 2, at the data selection stage, each actor performs several actions to update its memory and inform decision-making. Take domain actor as an example, it takes three-step action during data selection stage: (1) Sample pretraining data points from the pretraining data pool, distributing them uniformly across each domain; (2) Call the current model to assess the reward of each data point and gather feedback; (3) Update the memory of domain weights based on gathered feedback and adjusts the score for each data point by incorporating prior knowledge from the offline labeling process. This process is similarly followed by the quality actor and the topic actor.
Actor update.
After sampling data uniformly from actor search space, each actor updates its internal weights using local information based on the sampled data. For example, for domain actor, it calculates the average influence of each domain. For actor $\mathcal{A}\in\{\mathcal{A}_{\text{Quality}},\mathcal{A}_{\text{Domain}},\mathcal{A}_{\text{Topic}}\}$, it updates its internal weights by calculating the overall rewards sampled from each interval as:
$\vspace{-0.5em}\mathcal{R}(\mathcal{D}_{\mathcal{A}}^{j})=\frac{1}{|D_{\mathcal{A}}^{j}|}\sum_{x_{i}\in D_{\mathcal{A}}^{j}}r(x_{i})\coloneqq\overline{R}_{\mathcal{A}}^{j},\vspace{-0.5em}$ (4)
where $D_{\mathcal{A}}^{j}$ represents the sample set of the $j$-th subcategory under actor $\mathcal{A}$, e.g. Wikipedia for domain actor. And $x_{i}$ is a sample within this sample set. The sliding averaging is used to update the weight for each subcategory $w_{\mathcal{A}}^{j}$ with current rewards:
$w_{\mathcal{A}}^{j}\leftarrow(1-\eta_{A})\cdot w_{\mathcal{A}}^{j}+\eta_{A}\cdot\overline{R}_{\mathcal{A}}^{j},\vspace{-0.5em}$ (5)
where $\eta_{\mathcal{A}}$ is the sliding average factor to tradesoff bias-variance. The overall updated weight of actor $\mathcal{A}$ is an vector in $n_{\mathcal{A}}$ dimension, $\mathbf{w}_{\mathcal{A}}=[w_{\mathcal{A}}^{1},...,w_{\mathcal{A}}^{n_{\mathcal{A}}}]$, where $n_{\mathcal{A}}$ is the number of total subcategory within the space of actor $\mathcal{A}$. Utilizing the prior memory stored by each actor, it can give out a final score for each data point as $\mathcal{S}_{\mathcal{A}}(x_{i})=w_{\mathcal{A}}^{j}$, where $j$ is the subcategory that $x_{i}$ belongs to.
3.4 Multi-Actor Collaboration
Ultimately, the actor console defined in Definition 2 aggregates all actors’ feedback to compute a final score for each data point, determining the final data selection decision.
Multi-actor collaboration.
In the context of multi-actor collaboration, the weighted score for each actor must be calculated to evaluate their respective contributions effectively. This calculation takes into account various factors specific to each actor. For every data sample $x_{i}$, the overall score $S(x_{i})$ is determined by the following formula:
$S(x_{i})=(\theta_{\text{Quality}}\cdot\mathcal{S}_{\text{Quality}}(x_{i})+$ (6)
$\theta_{\text{Domain}}\cdot\mathcal{S}_{\text{Domain}}(x_{i})\ +\theta_{\text{Topic}}\cdot\mathcal{S}_{\text{Topic}}(x_{i})),$
where $\mathcal{S}_{\text{Quality}}(x_{i})$, $\mathcal{S}_{\text{Domain}}(x_{i})$, and $\mathcal{S}_{\text{Topic}}(x_{i})$ are scores calculated by quality, domain, and topic actors for sample $x_{i}$. $\mathbf{\theta_{\mathcal{A}}}\in\{\theta_{\text{Quality}},\theta_{\text{Domain}},\theta_{\text{Topic}}\}$ is the collaborative weight for each actor, which is updated during training process.
Table 1: Our approach boosts model performance across tasks. To fit demonstrations within 1024 tokens, we provide full results for 0, 3, and 5 shots in Appendix A.8. The table covers Problem Solving, Commonsense Reasoning, and Reading Comprehension, with best QuRating and DSIR variants: QuRating-Edu and DSIR-Wiki.
Selection Method Problem Solving Commonsense Reasoning Reading    Comprehension Average (4 tasks) (4 tasks) (2 tasks) (10 tasks) Random sampling - 30B tokens 31.1 32.9 43.1 34.2 Random sampling - 60B tokens 33.6$\uparrow$2.5 33.7$\uparrow$0.8 46.1$\uparrow$3.0 36.1$\uparrow$1.9 Perplexity PPL (Ankner et al., 2024) 29.9$\downarrow$1.2 30.5$\downarrow$2.4 42.4$\downarrow$0.7 32.7$\downarrow$1.5 Classifier-based data selection QuRating (Wettig et al., 2024) 34.1$\uparrow$3.0 34.1$\uparrow$1.2 41.4$\downarrow$1.7 35.6$\uparrow$1.4 FineWeb-Edu (Penedo et al., 2024) 32.6$\uparrow$1.5 33.0$\uparrow$0.1 45.3$\uparrow$2.2 35.3$\uparrow$1.1 DSIR (Xie et al., 2023b) 30.9$\downarrow$0.2 32.0$\downarrow$0.8 41.5$\downarrow$1.6 33.5$\downarrow$0.7 Domain mixing methods DOGE Fan et al. (2024) 30.9$\downarrow$0.2 32.2$\downarrow$0.7 45.1$\uparrow$2.0 34.3$\uparrow$0.1 DoReMi (Xie et al., 2023a) 30.4$\downarrow$0.7 32.6$\downarrow$0.3 44.8$\uparrow$1.7 34.1$\downarrow$0.1 DMLaw (Ye et al., 2024) 30.2$\downarrow$0.9 32.1$\downarrow$0.9 45.1$\uparrow$2.0 33.9$\downarrow$0.3 RegMix (Liu et al., 2024) 30.7$\downarrow$0.4 32.5$\downarrow$0.4 44.6$\uparrow$1.5 34.2$\uparrow$0.0 Influence MATES (Yu et al., 2024) 30.9$\downarrow$0.2 34.0$\uparrow$1.1 46.5$\uparrow$3.4 35.3$\uparrow$1.1 Multi-actor collaboration (ours) 36.7$\uparrow$5.6 34.8$\uparrow$1.9 45.9$\uparrow$2.8 37.8$\uparrow$3.6
Collaborative weight update.
To dynamically adjust the importance of each actor during various training phases, we modify the actor’s collaborative weight based on its overall rewards. We compute the reward of each actor and the average reward across all actors:
$\overline{R}_{\mathcal{A}}=\frac{1}{|n|}\sum_{j=1}^{n}w_{\mathcal{A}}^{j}\cdot\overline{R}_{\mathcal{A}}^{j},\quad\overline{R}=\frac{1}{3}\sum_{\mathcal{A}}\overline{R}_{\mathcal{A}},$ (7)
This information is then used to update each actor’s collaborative weight, which is stored in the actor console’s memory for future decision-making:
By continuously refining these weights, the collaboration strategy adapts to optimize overall performance and appropriately adjust the role of each actor throughout different stages of training. Complete training pipeline is outlined in Algorithm 2.
4 Experiments
We conduct a series of experiments to evaluate the effectiveness of our multi-actor collaborative data selection method. Comprehensively, we find that: (1) In the end-to-end experiments, our approach introduces significant improvement in terms of data efficiency leading to faster convergence for LM training, and achieves up to $10.5\%$ improvements on average across various language model benchmarks when compared with other baseline approaches (§4.1); (2) We also verify that the design and implementation of the core components in our multi-actor framework design are necessary to reach this advanced performance through a set of carefully designed ablation studies (§4.2).
4.1 End-to-end Experiments
We evaluate our multi-actor framework against a wide category of state-of-the-art approaches to compare the data efficiency for LM pretraining. We train a $1.3$ billion parameter Llama-2 architecture model with $30$ billion selected tokens. We also report our results of generalizing our methods to 3.6B and 8B models in Appendix A.1.
  • Pretraining datasets. We utilize the popular SlimPajama (Soboleva et al., 2023) dataset including $627$ billion tokens, which is derived from the RedPajama (Computer, 2023) dataset. The SlimPajama (Soboleva et al., 2023) provide the meta-data about the domain information for each sample. Before the training process, we annotate the entire dataset using the FineWeb-Edu quality scorer (Penedo et al., 2024) along with our custom-trained BERT-based topic classifier. The training details for the topic classifier is provided in Appendix §A.2.
Table 2: This ablation study examines the performance of various combinations of actor collaboration and update mechanisms. All models are in 1.3B LLaMA2 architecture. Three-shot accuracy is reported for all tasks, with the highest value in each column shown in bold.
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Quality&Domain&Topic actor 65.8 31.5 23.0 26.6 24.6 39.9 54.1 20.1 60.4 30.5 37.7 without collaboration update 59.4 26.3 21.3 25.1 20.5 38.9 52.9 19.8 58.1 28.3 35.1 Domain&Quality actor 63.3 29.7 22.6 25.1 21.8 40.5 53.1 20.3 59.5 28.8 36.5 Topic&Quality actor 62.9 28.1 22.3 26.5 22.6 39.6 51.8 21.7 56.7 30.7 36.3 Domain&Topic actor 55.6 25.2 21.8 26.5 23.1 39.1 53.7 20.9 57.5 29.0 35.2 Quality actor 59.1 29.7 22.4 25.3 21.1 38.5 51.2 19.1 57.2 28.3 35.2 Domain actor 54.1 25.6 21.4 25.9 22.3 38.1 53.6 20.0 58.1 27.9 34.7 Topic actor 55.3 25.3 21.9 27.1 22.1 39.4 51.5 19.8 56.3 28.9 34.8 No actor 54.6 23.0 22.1 24.9 18.8 40.3 52.9 21.5 53.0 29.8 34.1
Discussion. We highlight that our proposed multi-actor collaborative data selection mechanism introduces statistical efficiency in terms of LM training convergence and also provides some computational efficiency in terms of data processing overheads. In terms of statistical efficiency, our method consistently outperforms others at every benchmarked training step, as shown in Figure 3. While MATES Yu et al. (2024) performs comparably to our methods during the early training phase (steps $1500$ to $3000$), its performance drops in later stages. This aligns with its original paper, which notes that relying solely on influence functions for specific reference tasks (e.g., LAMBADA (Paperno et al., 2016)) can degrade performance in mid-to-late pretraining. Despite this, MATES still outperforms other methods without dynamic adjustments shown in Figure 3. In contrast, our multi-actor collaborative data selection mechanism can dynamically adjust the corresponding weights from different actors and select data based on the most up-to-date model preferences, effectively mitigating biases and surpassing other domain-mixing and data-selection techniques. In terms of computational efficiency, we also achieve higher computational efficiency than previous methods. For example, QuRating Wettig et al. (2024) requires around $7.13\times 10^{20}$ FLOPs to label the entire SlimPajama dataset, while our offline labeling takes just $9.91\times 10^{19}$ FLOPs. MATES Yu et al. (2024), which recalculates influence scores and trains a BERT model for each labeling cycle, incurs $1.98\times 10^{20}$ FLOPs for a four-stage update. Additionally, MATES’ labels are only usable in the next training stage, making it time-consuming and difficult to scale. In contrast, our method can improve the computational efficiency from two aspects: (1) we find that a group of light-weight actors collaboratively enables superior data selection, which is more computational efficiency than any method that requires a heavy data processing or label procedure; (2) the collaborative, dynamic learning procedure introduced in our multi-actor framework is computational efficient; by using a sampled holdout set and CPU-based calculations for updating actor parameters, our computational overhead is ignorable compared with heavy LM training computation.
4.2 Ablation Study
We introduce a set of carefully designed ablation studies to justify the design and implementation of our multi-actor collaborative data selection framework. Concretely, (1) we test the combination of different actors to show the advance introduced by collaboration, and (2) we verify the necessity of the dynamic adjustment of the actor’s weight for data selection.
Results and discussion. The results of the ablation study are shown in Table 2. We want to highlight the result from two aspects: First, the ablation study underscores the importance of each actor in achieving optimal performance across the training tasks. When the quality, domain, and topic actors are used together, the model performs best, highlighting the benefits of their combined use, as shown in Table 2. In terms of evaluating each actors’ contributions, we find that the quality actor excels in problem-solving tasks like ARC-E and MathQA by leveraging educational knowledge but is less effective for domain-specific or context-heavy tasks like BoolQ and RACE. The domain actor enhances commonsense reasoning (e.g., CommonsenseQA) and reading comprehension (e.g., BoolQ) by incorporating domain-specific knowledge. The topic actor is most effective for multi-topic tasks like MMLU and contributes significantly to commonsense reasoning tasks like SocialIQA. Second, the ablation study verifies the design and implementation of the collaborative dynamic adjustment of the actors’ weights (introduced §3.4) for efficient data selection. When actors were initialized with equal, fixed weights instead of using dynamic weighting, overall performance dropped significantly, as shown in Table 2.
5 Related Work
Data selection in LM pretraining. Selecting high-quality pretraining data from large corpora is crucial for effective LM training. Recent approaches leverage various methodologies for efficient data selection. Concretely, classifiers (Brown et al., 2020; Chowdhery et al., 2023; Du et al., 2022; Xie et al., 2023b) and language modeling perplexity (Wenzek et al., 2020; Thrush et al., 2024) have been applied to identify data resembling high-quality samples; recently, more advanced quality scores based on classifier have shown the effectiveness in data selection, e.g., QuRating (Wettig et al., 2024), FineWeb-Edu (Lozhkov et al., 2024), etc. Data mixture is another effective way to improve data diversity, at both token level (Touvron et al., 2023a; Gao et al., 2020; Soboleva et al., 2023) and sample level, e.g., DoReMi (Xie et al., 2023a), DOGE (Fan et al., 2024), DMLaw (Ye et al., 2024), and RegMix (Liu et al., 2024); very recently, topic distributions has also been considered as an effective data mixing method, e.g., the downsampling overrepresented topics in Llama 3.1 (Team, 2024). Influence functions have been studied to understand for data efficiency  (Koh and Liang, 2017), and some recent attempts based on efficient approximation have been proposed to improve data efficiency in LM pretraining (Schioppa et al., 2022; Grosse et al., 2023; Isonuma and Titov, 2024); For example, MATES (Yu et al., 2024) uses a staged BERT model to assess data influence, QUAD (Zhang et al., 2024) leverages cluster information to reduce the computational cost of calculating individual data influence.
Multi-agent and multi-actor collaborative frameworks. Collaboration across multiple autonomous or semi-autonomous entities has been studied extensively under the paradigms of multi-agent and multi-actor systems. Traditional multi-agent systems (Russell and Norvig, 2016; Wooldridge, 2009) involve autonomous agents that make independent decisions and learn through interaction, often coordinated via mechanisms such as reward signals, negotiation, or centralized planning. These systems have been effective in applications such as neural architecture search (Bello et al., 2017), collaborative large language model programming (Hong et al., 2024), and distributed control (Olfati-Saber, 2006). In contrast, our work adopts a multi-actor perspective, where actors are not general-purpose intelligent agents, but specialized components—specifically, independently operating data selection methods. Each actor follows a distinct heuristic to prioritize pretraining data, adapting dynamically to the model’s evolving state. A central coordination mechanism integrates these priorities over time, resolving conflicts and guiding collective behavior. This aligns with formalizations of multiactor systems, where the term "actor" serves as a generic abstraction that encompasses agents, bodies, or effectors (Russell and Norvig, 2016). Unlike multi-agent systems, which may involve deliberative autonomy and decentralized goal negotiation, multiactor systems often assume a shared objective and focus on concurrent action, coordination, and conflict resolution. Our framework draws from both traditions: it retains the modularity and adaptability of multi-agent systems, while leveraging the structured concurrency and coordination principles emphasized in multiactor planning (Ligtenberg et al., 2001, 2004; Wang et al., 2020). This hybrid approach enables us to tackle a largely underexplored challenge: resolving competition among heterogeneous data selection heuristics to improve the efficiency and effectiveness of language model pretraining.
6 Conclusions
In this paper, we introduce a multi-actor collaborative data selection framework to enhance efficiency in LM pretraining. Our framework allows multiple data selection methods to operate as independent actors, with an actor console designed to dynamically integrate their outputs throughout the LM training process. Empirical studies show it improves data efficiency, speeds up convergence, and achieves up to 10.5% average performance gains over state-of-the-art methods. These results demonstrate the effectiveness of dynamically combining data selection strategies to resolve conflicts and optimize LM pretraining.
Limitations
While our method greatly improves data selection for language model pre-training, our study has some limitations. Due to computational constraints, our experiments were limited to relatively small-scale models (up to 8B parameters) with restricted token budgets. Additionally, while our quality metrics are comprehensive, they may not fully capture all dimensions of pre-training data. In future work, we plan to refine or expand these metrics to bridge these gaps.
Acknowledgement
This work is supported by the HKUST startup grant R9895 from CSE, RGC-ECS project 26218024, Shanghai Artificial Intelligence Laboratory, the National Key R&D Program of China (2022ZD0160201), the National Key R&D Program of China (2024YFA1014003), National Natural Science Foundation of China (92470121, 62402016), and CAAI-Ant Group Research Fund. We express sincere thanks to InternTrain Team of Shanghai Artificial Intelligence Laboratory, especially Yang Gao, for their kind help for the pre-training experiments.
Appendix A Appendix
Table 3: This study compares the performance of training 3B model from scratch on 60B tokens selected using random sampling versus multi-actor collaboration. Accuracy is reported for all tasks, with the highest value in each column shown in bold.
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random Sampling 34.8 17.7 21.3 23.0 12.0 32.9 50.2 19.6 37.8 20.9 27.0 Multi-actor collaboration (Ours) 42.9 21.3 21.9 24.0 15.8 33.9 51.0 20.4 54.8 21.2 30.7
Table 4: This study compares the performance of training 8B model from scratch on 25B tokens selected using random sampling versus multi-actor collaboration. Accuracy is reported for all tasks, with the highest value in each column shown in bold.
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random Sampling 22.2 53.3 21.5 23.3 19.4 36.1 51.0 18.2 48.0 21.3 31.4 Multi-actor collaboration (Ours) 25.5 58.0 23.2 24.1 21.6 38.8 53.1 20.8 57.8 22.9 34.6
A.1 Generalization to 3.6B and 8B Models
To evaluate the scalability of our approach, we conducted additional experiments training a 3.6 billion parameter model based on the LLaMA 3.2 architecture, which further demonstrates the scalability of our method. So far we have trained on 36 billion tokens and achieved strong performance, with plans to continue training with additional tokens according to scaling laws. As Table 3 shows, when compared to random selection, our method shows consistent performance improvements across all downstream tasks, achieving a 13.7% increase in average accuracy—significantly higher than the 10.5% improvement observed with the 1.3B models. Our results in Table 4 show that our methods outperform random selection approaches by 10.2% for 8B LLaMA 3.2 architecture models. Based on trends across three different model sizes (373M, 1.3B, 3.6B and 8B), our approach consistently outperforms random selection by over 10% on average. This consistent advantage makes us believe that it suggests our method has strong potential for training even larger models, including those with 10B+ parameters.
A.2 Details of Training Topic Classifier
@@
A.3 Guidelines for Generalizing a New Criteria as an actor
This section provides a detailed guidelines for incorporating a new criterion into our multi-actor system. The process is designed to ensure seamless integration and effective collaboration between existing and new actors.
Algorithm 1 Integrating a New Criterion into Multi-Actor Collaboration
1:Sampled dataset $\mathcal{D}_{\text{sample}}$, pretraining dataset $\mathcal{D}_{\text{train}}$, reference dataset $D_{\text{ref}}$, existing Actors $\{\mathcal{A}_{i}\}_{i=1}^{3}$, scoring weights $\{\theta_{i}\}_{i=1}^{|\mathcal{D}|}$, memory $\mathcal{H}_{\mathcal{A}}$ for each Actor.
2:Annotating Data for the New Criteria
3:    Sample data from the whole datasets, and annotate sampled dataset $\mathcal{D}_{\text{sample}}$ according to new criterion.
4:Training a Classifier for the New Criteria
5:    Train a supervised classifier on $\mathcal{D}_{\text{sample}}$.
6:Defining the New Actor
  • Action Space: Sample and assess data points across subcategories of the new Criteria.
  • Memory: Store prior scores and update based on model feedbacks.
7:Labeling the Pretraining Dataset
8:    Use trained classifier to label the whole training dataset $\mathcal{D}_{\text{train}}$ and store these labels in memory $\mathcal{H}_{\mathcal{A}_{\text{new}}}$.
9:Defining Actor Weights and Collaboration Strategy
  1. 1.
    Define the subcategory weights updating mechanism for $\mathcal{A}_{\text{new}}$ using Eq. 5:
    $w_{\text{new}}^{j}\leftarrow(1-\eta_{\text{new}})\cdot w_{\text{new}}^{j}+\eta_{\text{new}}\cdot R_{\text{new}}^{j}.$
  2. 2.
    Integrate $\mathcal{A}_{\text{new}}$ into the scoring function using Eq. LABEL:eq:coordinator:
    $S(x_{i})=\theta_{\text{Quality}}S_{\text{Quality}}(x_{i})+\theta_{\text{Domain}}S_{\text{Domain}}(x_{i})+\theta_{\text{Topic}}S_{\text{Topic}}(x_{i})+\theta_{\text{new}}S_{\text{new}}(x_{i}).$
10:Actor Initialization with Regression Techniques
11:    Initialize the Actor weights $w_{\text{new}}$ using regression techniques (e.g., RegMix).
Our framework offers several significant benefits when integrating a new actor. First, it offers flexibility, as the addition of new criteria can be performed independently of the core framework. This decoupling ensures that introducing new components does not require significant structural changes, allowing for smooth integration with minimal disruption to existing processes. Second, the approach is highly scalable. By enabling the training of new classifiers offline, the system can be easily adapted to handle a wide variety of data selection goals, which can evolve over time as new criteria emerge. Finally, the framework ensures extensibility, meaning it can seamlessly accommodate new objectives, whether simple or complex. This extensibility is key to maintaining efficient and effective collaboration across multiple actors, regardless of the size or complexity of the task at hand.
As demonstrated in Algorithm 2, our framework seamlessly integrates a new actor through a series of straightforward steps. This approach ensures that the multi-actor framework remains flexible and effective in addressing diverse data selection objectives while preserving its collaborative efficiency.
A.4 Connection of Multi-actor Collaborative Selection Method to Multi-agent RL
The proposed multi-actor collaborative selection method is fundamentally inspired by the intelligent actor defined in Russell and Norvig (2016), where the actor generally refers to an entity that perceives some status and map the observed status into actions. However, our framework also has many similarity compare with traditional multi-agent framework in reinforcement learning, where multiple actors work together to optimize a shared objective. In this section, we formally demonstrate the relationship between our framework and traditional multi-agent RL.
A.4.1 Overall Definition in Reinforcement Learning Formulation
We first clearly formulate each component of framework compare with components in general MARL framework for understanding the mechanism of our framework. As our goal is to select the opt the global action at each step involves selecting a subset of data, $\mathcal{D}_{k}$, from the entire dataset, $\mathcal{D}$. This subset is used to update the model, where batches are drawn from $\mathcal{D}_{k}$ for training. The global state is represented by the current model parameters, $M$, which evolve as the model is trained. The state transition is formalized as $M^{\prime}=\mathcal{O}(M,\mathcal{D}_{k})$, where $M^{\prime}$ denotes the updated model after training on $\mathcal{D}_{k}$.
The reward function measures the improvement in model performance on a reference task, $\mathcal{T}_{\text{ref}}$, which serves as a proxy for the true downstream task $\mathcal{T}_{\text{eval}}$. The reward is defined as:
$R(M^{\prime}|M,\mathcal{T}_{\text{ref}})=\mathbb{E}_{x\sim\mathcal{T}_{\text{ref}}}[-l(x;M^{\prime})],$
where $l(x;M^{\prime})$ is the loss on $\mathcal{T}_{\text{ref}}$. For individual data points, the reward is estimated using influence functions:
$r(x_{i})=\text{Influence}(x_{i},M,\mathcal{T}_{\text{ref}}).$
This formulation links data selection directly to its impact on improving the model’s performance on $\mathcal{T}_{\text{ref}}$.
A.4.2 Actor Design
The estimation of value is as follows: The actor stores a reward estimation vector for each subset. The update rule is given by
$w_{\mathcal{A}}^{j}(t+1)=(1-\eta_{\mathcal{A}})\cdot w_{\mathcal{A}}^{j}(t)+\eta_{\mathcal{A}}\cdot\bar{R}_{\mathcal{A}}^{j}.$
The sliding average is used here because if all data in a subset were fully processed to compute $\bar{R}_{\mathcal{A}}^{j}$, there would be no need for a sliding average. However, since only a portion of the data is sampled, the estimate has higher variance, which is not favorable for training. At the same time, the influence score itself is dynamic (even if the data remains constant, the model evolves). Averaging with outdated scores introduces bias. Therefore, the sliding average factor $\eta_{\mathcal{A}}$ strikes a ’bias-variance tradeoff’. We assume the score estimate for each data point $x_{i}$ in $\mathcal{D}_{\mathcal{A}}^{j}$ with respect to the dimension of interest for the actor, is given by $S_{\mathcal{A}}(x_{i})=w_{\mathcal{A}}^{j}$ where $x_{i}\in\mathcal{D}_{\mathcal{A}}^{j}.$
Algorithm 2 Multi-actor collaborative data selection for LM pretraining